OpenStudy (anonymous):

Insert 4 Gm.s between 2 and 486

ganeshie8 (ganeshie8):

2, _, _, _, _, _ 486

ganeshie8 (ganeshie8):

the question is simply asking us to construct a geometric sequence with : first term = 2, sixh term = 486

OpenStudy (anonymous):

Yes , is the common ration of any GP = \[\huge [\frac{ b }{ a }]^{\frac{ 1 }{ n+1 }}\]

ganeshie8 (ganeshie8):

i don't like memorizing a formula for such a simple thing, but the formula would be : \[\huge [\frac{ b }{ a }]^{\frac{ 1 }{ n\color{Red}{-}1 }}\]

OpenStudy (anonymous):

in book it is given n+1

ganeshie8 (ganeshie8):

lets derive it :)

OpenStudy (anonymous):

Okay...

ganeshie8 (ganeshie8):

we're given below : first term = a nth term = b

ganeshie8 (ganeshie8):

nth term of geometric sequence is given by : \[\large b = ar^{n-1}\]

OpenStudy (anonymous):

oh got it

ganeshie8 (ganeshie8):

isolate \(\large r\) ^^

ganeshie8 (ganeshie8):

once you have the value of \(r\), you can find the 2nd,3rd,4th and 5th terms

OpenStudy (anonymous):

so it is just manipulating terms , but the answer is 6 ,18 , 54 ,162

ganeshie8 (ganeshie8):

thats right !

ganeshie8 (ganeshie8):

first term = 2 6th term = 486 so, \[\large 486 =2*r^{6-1}\]

ganeshie8 (ganeshie8):

\[\large 243 =r^{5}\]

ganeshie8 (ganeshie8):

\[\large r=3\]

ganeshie8 (ganeshie8):

multiply the first term by \(3\) to get second term multiply the second term by \(3\) to get third term and so on...

OpenStudy (anonymous):